if x+1/x=5 wat is x-1/x

Answer by waynest(281)

(x+1)/x=5
x(x+1│x)=(5)x
x+1=5x
1=5x-x
1=4x
x=1/4
If x=1/4 then
(x-1)/x=
(1/4-1│1/4)=
-(3/4)/(1/4)=
-3/4×4=-3